# Search and Society Reimagining Information Access for Radical Futures

### v1

@@ 2 Background
Transparency in IR covers a broad range of scenarios and concerns. Examples include transparency about how specific components of the system behave (e.g., query reformulation [270] and ranking [328, 364, 327, 409, 75, 77, 326, 366, 169]) and how data subjects are represented in search results (e.g.[48, 226]). Transparency needs may specifically arise in the context of how information and knowledge access systems modulate what and who get exposure and influence how we see ourselves and others [84]. Different notions of transparency may be relevant here, including but not limited to: System transparency (i.e., how does the system work?), procedural transparency (i.e., in what social norms and processes is the system use embedded?), or transparency of outcomes (i.e., what is the impact of the system’s use on individuals and society?). New transparency needs [231] may also arise in the context of emerging technologies, such as large language models (LLMs) [278, 349, 78, 351]. However, an important question that all transparency research must contend with is: transparency towards what end? Some works (e.g.[293, 314]) motivate transparency as a means to increase user trust in the system. However, we should be critical of whether that trust is warranted, or whether transparency mechanisms could in fact draw users into a false sense of safety and distract them from noticing how the system surveils them and subtly manipulates their behavior [302, 269]. Indeed, Hollanek [175] argue that “only the sort of transparency that arises from critique—a method of theoretical examination that, by revealing pre-existing power structures, aims to challenge them—can help us produce technological systems that are less deceptive and more just”. It is that kind of critical reflection that we believe should inform transparency research in IR and lead the community towards explicit goals to challenge power (e.g., the power that system owners hold over users) and safeguard user agency.
@@
On the same lines, misinformation research should be motivated by sociotechnical visions for the future of democratic societies, public health, and knowledge production. When we broaden that frame, it becomes apparent that the community must not predominantly focus on automated fact checking, also an important research problem, but understand the social, political, and economic conditions under which misinformation and disinformation is produced and disseminated. The focus of misinformation research then should include identifying, understanding, and addressing the structural mechanisms of misinformation (e.g., data voids [153]) as well as ground itself in the articulation of IR’s role in online knowledge production, public health education, and information literacy.
@@
Our perspective here on ongoing research on fairness, transparency, and ethics in IR should not be misconstrued as an argument for doing less of this kind of work. Instead, we believe that the community should be explicit and more ambitious about the changes it wants to affect in broader society and conduct research with a clear mapping between the research goals and the desired social impact. Just as Johnson [195] challenge the notion that open data directly leads to information justice, we want the IR community to be cautious in their assumption that working on narrowly defined questions of fairness, transparency, and ethics necessarily contributes towards practical social good. To be effective in that endeavour, we believe that we should be explicit in articulating our collective visions for our sociotechnical futures, the changes we want to affect in society, and how we envision our research can bring about those changes.
3 Why now?
This is an exciting time for IR research. Research communities constantly evolve, shaped by ideas and developments from both within the field and adjacent communities, and in response to real world events and changing societal needs. In case of IR, we believe we are seeing significant developments in both context: (i) A fast-changing landscape across IR and adjacent fields, such as natural language processing (NLP), human-computer interaction (HCI), machine learning (ML), and artificial intelligence (AI), spurred by recent progress in LLMs, and (ii) an increasing recognition of the role of technology, and the communities that build it, in determining our collective futures.
Consequently, there is a shared sense in the community that right now both IR technologies and IR research are undergoing transformative changes [27]. This presents a timely opportunity for the field to consciously, collectively, and ambitiously engage in purposeful dialog about the future of IR. In doing so, we must also critically reflect on “where do we want to go?” (i.e., our sociotechnical imaginaries), “how do we get there?” (i.e., our theories of change), and “who will we go there with?” (i.e., our relationships, and that of our work, with other disciplines, industry, and society). These considerations should drive future IR research as a whole, and we should accept this opportunity to re-center our research agendas on societal needs while dismantling the artificial separation between the work on fairness and ethics in IR and the rest of IR research.
LLMs are changing how we access information. The natural language generation capabilities of LLMs are having a profound effect on how we access information and in what context. Conversational search interfaces have gone from being aspirational [17, 257] to being deployed at web-scale (e.g., Bing Chat [^3] https://chat.bing.com/ (now Microsoft Copilot [250]) and Google Bard [^4] https://bard.google.com/) in a span of two years. While the long-term social implications of inserting an LLM between a retrieval system and the information seeker should rightly be met with rigorous skepticism [317, 365], natural language interfaces are already impacting how we interact with IR systems. The ubiquitous search box is also being challenged as IR becomes more context-driven than user-driven as a consequence of LLMs increasingly embedding themselves in user’s work processes (e.g., Microsoft Copilot for M365 [251, 376]) and interacting with the IR system on the user’s behalf, under retrieval-augmentation [225, 400].
While we should be excited with the new prospects that these emerging AI technologies unlock and recognize that they will shape how we access and interact with information in the future, we must not be duped by AI techno-determinism into believing that there is a single pre-determined path forward. Instead, we must hold pluralistic views of what IR’s future, one that is yet to be determined, looks like and how these technologies will take us there. In a study of top-cited AI papers, many of which are coauthored by researchers affiliated with industry or elite universities, Birhane et al. [58] find that the dominant values expressed and operationalized support concentration of power. So, we must ask: in what new ways can we imagine accessing and interacting with information, aided by LLMs, if large-scale IR systems were not just a purview of big tech? How would LLMs empower us to reimagine IR systems whose explicit goal is to dismantle hierarchies and redistribute power, not to centralize it? What role would AI technologies play in information access that is built explicitly to facilitate dialogical social processes for knowledge production, world building, and our collective struggles for universal emanicipation? It is critical that we have these conversations now in the face of ongoing massive technology-driven power shifts in favor of dominant established platforms that grants their visions of the future normative status and shrinks the space for any critique, resitence, or counter-imaginaries.
LLMs are shifting priorities of IR research. Over the last decade, deep learning technologies became the new hammer in the toolbox for IR research [263, 233, 122], dominating IR publications with nearly four out of five papers at the ACM SIGIR 2020 conference being related to deep learning by some estimates [262]. One particular focus of neural IR has been on estimating relevance of information artefacts (e.g., documents) to an information intent (e.g., as expressed by a search query) for ranking, a central problem tin IR. Curiously, many of the key ingredients for this research, such as the Transformer architecture [361] and the idea of pretrained LLMs, like BERT [98], came from fields adjacent to IR; correspondingly, shifting the focus within the field more towards adapting these models for the relevance estimation task (e.g., [275]) and making them more efficient [69].
More recently, Thomas et al. [347] demonstrated that LLMs, like GPT-4 [278], are able to estimate the actual searcher’s preference for documents, given their query, better than several populations of human relevance assessors. This technology has already been deployed in production at Bing. [^5] https://twitter.com/IR_oldie/status/1659413086007328768 Putting it bluntly, these LLMs may be getting close to the best we can expect with machine learned general purpose relevance estimators. If these claims stand the test of time, it may mark a watershed moment for IR research. Speculatively, we may see the IR community further shifting towards improving efficiency of these models, focusing on more challenging information needs (e.g., tip-of-the-tongue queries [21, 22]), and increasing investments in measurement and evaluation (e.g., for emerging new IR scenarios, such as retrieval-augmented generation). Alternatively, we may ask: How can the IR community meet this moment, not with apprehension nor with unchallenged exuberance for progress happening in adjacent fields, but truly grasp this opportunity to redefine what it means to work on IR research? Can we be unabashedly discontent with imagining the future of IR based wholly on what new AI progress makes plausible, and instead reimagine our field as a place where knowledge, culture, and radical aspirations meet to demand of technology to make new futures possible? Alternatively, if we fail to articulate an aspiring vision for IR research, we risk as a field being reduced to just an application of AI.
LLMs are raising new sociotechnical concerns. It is well-known that language models reproduce, and even amplify, harmful stereotypes and biases of moral import [137] that are present in their training data [60, 73, 154, 59, 41, 3]. One particular mitigation strategy involves using ML approaches that learn from human preferences, such as reinforcement learning from human feedback (RLHF) [79, 410], to align with “human values” [204, 342]. While RLHF has been quite effective in constraining LLMs from producing certain types of offensive and harmful content, we must be wary of any framing of AI ethics, such as AI alignment [308, 138, 139], that presupposes the existence of universal values [295, 57, 194, 291, 310]. This is particularly concerning if we look at this in the context of power asymmetries that exist between powerful private corporations, who have outsized influence over what values these models are optimized for, and those who use these models or are represented in some fashion in the model outputs. This is further compounded by the lack of appropriate mechanisms for civil society to participate in and challenge these choices. Indeed, by placing these controls in the hands of the privileged few, we risk further concentration of power. The concerns of biases in what these models produce, and even biases in what context they refuse to generate [357], and who gets to influence those decisions have serious implications for information access and society. Finally, the development of LLMs themselves may involve potential harms to authors [96, 221, 68, 321, 261], crowdworkers [158, 114, 159, 305, 197, 306, 384, 285, 112], and even the environment [41, 283, 61, 388, 105, 284].
In a critical perspective, Shah and Bender [317] recommend that IR research should focus on developing appropriate guardrails in anticipation of the social implications of these emerging technologies but not be constrained by a singular LLM-powered conversational search vision for IR. With this we agree, we should be excited by the new capabilities unlocked by recent progress in LLMs but must not limit our imaginations and aspirations by only what LLMs make plausible.
Our relationships with adjacent communities are changing and so is how we do research. Not long ago, many in the IR community would tout IR, and specifically web search, as a rare success story of real-world application of ML, AI, and NLP technologies. A sea change in these adjacent communities in the last decade have shifted this balance. Now, many see IR as just another NLP task, sometimes included in NLP benchmarks (e.g., HELM [229]) for evaluating ML/AI models. With retrieval-augmented LLMs, IR is auditioning for a new role, as a tool for AI models, curiously inverting the relationship between these technologies, where AI was one of many in the IR toolbox.
These communities are also undergoing significant changes in research culture, often influencing each other on the way. One particular trend in NLP, and the broader ML and AI communities, that has influenced IR, is leaderboard-driven research. Several NLP leaderboards (e.g., [29, 157, 198, 301, 373, 341, 374, 220, 396, 230]) have been instrumental in encouraging progress on specific tasks. IR has a long history of focus on shared tasks and benchmarks, notably TREC [369] that has been a venue for developing new tasks and benchmarks, as well as building research communities with shared interests around them. What differentiates IR benchmarking in venues such as TREC from NLP leaderboards is that the former is framed not as a competition, but as a coopetition. In a competition, the goal of the participant is to outperform others, while in a coopetition the participants share a collective goal to develop a better understand of both the task and the models through robust evaluation. In the words of Ian Soboroff: [^6] https://twitter.com/ian_soboroff/status/1426901262369439751 “The datasets were not built to be solved. They were built as tools to understand the problem and the systems we build to ‘solve’ them.”; or as Voorhees [368] put it: “Coopetition is defined as competitors cooperating for the common good… While competition can give one a bigger piece of the pie, cooperation makes the whole pie bigger.” By emphasizing the goals of community development and understanding of the tasks and the models, these evaluation effort try to promote scientific enquiry over sportive competition. Even the MS MARCO ranking task that initially started as a competition later reframed itself as a coopetitive evaluation effort [86, 235].
As a community we should cautiously embrace insights and trends from our neigboring fields. However, we should not let IR be minimized to just an evaluation task which undermines the critical responsibility that IR researchers owe to the broader society. Similarly, while leaderboards and competitions may be effective in creating excitement and increasing participation in certain tasks, we must be mindful of the implications of potentially a large section of the IR community being driven predominantly by these practices. When the goal is to win, then scientific enquiry takes a back seat, and the ones with the most compute and data resources take the metaphorical steering wheel. It risks, what Gausen et al. [143] call, albeit in a different context, shifting from praxis (“reflection and action directed at the structures to be transformed” [133]) to proxies (optimizing towards proxy quantitative measures of outcomes). Actions in this context may refer to any research activity, including but not limited to: formalization, design, experimentation, publishing, artefact creation, open sourcing, and community building; and examples of proxies include state-of-the-art (SOTA) performance on benchmarks and leaderboard rankings that do not translate to better scientific understanding or positive impact on people. [^7] This is a case of Goodhart’s law [80, 156, 180, 348] whereby improvements on benchmarks and corresponding metrics do not translate to progress on the problem the benchmark was created for, as has been argued for example by Hsia et al. [181].
Yet another relationship that we must critically examine is the one between industry and academia. Whittaker [379] point out that the concentration of data and compute resources, two key ingredients in recent advances in AI, in the hands of few large tech corporations is giving these same institutions tremendous power to shape academic research agenda. Big tech also shapes academic research agendas in various other ways, including academic engagements and employments. In IR, the MS MARCO dataset [29] and leaderboard [86, 234, 235], and the TREC Deep Learning track [85], that has been broadly adopted for benchmarking deep ranking models were produced and is currently maintained by industry researchers. Indeed, the organizers behind these efforts themselves recognized [86, 235] the critical responsibility that comes with defining critical research tasks for the community—effectively playing “the Pied Piper guiding a significant section of the community down specific lanes of research”—and recommend all benchmark developers to engage in open and inclusive discussions with the rest of the community to critically examine their impact. While academia-industry collaboration is critically important for the field to ground our research in real large-scale systems and see our research outputs materialize into real impact on system users, we must also resist the homogenization of our research agendas towards a singular world view put forth by big tech capitalism.
The world is changing and so is our relationship to that world. Our world at large is experiencing a confluence of many simultaneous, and mutually reinforcing, forces that are increasingly pushing us towards precarity, including but not limited to: rising global conflicts [345, 356], pandemic [315, 346, 129], and impending climate catastrophes [280, 184, 294]. At a moment when the world needs global solidarity built on trust and consensus, and informed citizenry with robust access to reliable information, online disinformation and misinformation are undermining both [355, 352, 205, 14, 106, 39, 399, 340]. While these complex global challenges require sophisticated and multifaceted response that spans across the political, legal, economic, and technological realms, one thing is for certain that information access research has a role to play. So, will we answer the call?
@@ 4 Towards sociotechnical change
Sociotechnical imaginaries are not born in vacuum. They are moulded and shaped by our values and our politics. Deliberation over what futures we want to bring into being is essentially political, and challenges us to critically reflect on our community’s shared, and yet pluralistic, political values. While the call for explicit political reflection in IR may come as a surprise to some, we need to recognize that our research and the artefacts we produce do not exist outside of the current sociopolitical order [136, 128, 260] but as essential cogs in the system, and the absence of political reflection does not imply an absence of politics in our work, but rather translates to implicit complicity in propping up the status quo and neoliberalism [107, 127, 208]. Instead, we can learn from how some of our neighboring fields, e.g., HCI and AI, have engaged with these questions, and reflect on how politics shapes and intersects with our own research agendas.
@@
There are several strands of research in IR-adjacent fields that explicate prefigurative politics [23] and ground research in humanistic [32, 31], anti-oppressive and emancipatory [330, 32, 202, 265, 312], feminist [371, 372, 33, 35, 37, 36, 34, 100], queer [232, 209, 165], postcolonial and decolonial [186, 287, 108, 338, 12, 13, 11, 187, 6, 264], anti-racist [2], anti-casteist [201, 310, 358, 359, 323, 203], anti-fascist [249], abolitionist [43, 30, 113, 196, 385], post-capitalistic [127, 66], and anarchist [208, 237, 24] epistemologies. Reviewing this full body of literature is out-of-scope of this work but we briefly present a sample to draw from and motivate new IR research agendas for sociotechnical change.
@@
Bardzell and Bardzell [32] define humanistic HCI as “any HCI research or practice that deploys humanistic epistemologies (e.g., theories and conceptual systems) and methodologies (e.g., critical analysis of designs, processes, and implementations; historical genealogies; conceptual analysis; emancipatory criticism) in service of HCI processes, theories, methods, agenda-setting, and practices”, and include emanicipatory HCI as an aspiration of humanistic HCI. Kane et al. [202] propose to incorporate emancipatory pedagogy [133] that does “not advocate the oppressed simply rise and overthrow their oppressors. Instead, […] the oppressors and oppressed create new educational processes that would allow them to work together to create a new type of society that was emancipatory for all”.
@@
In Science and technology studies (STS), there is a body of work [240, 370, 183, 354, 172, 367, 170, 259] on gendered inequities caused by technology, and how technology and gender relations mutually shape each other [371, 372]. In HCI, Bardzell et al. [33, 35, 37, 36, 34] propose to incorporate feminist theories [214, 135] into research and practice. Bardzell [33] posits that feminist theories can contribute to interaction design both by critiquing and by generating new insights that inform and shape designs and design processes. We can see feminist HCI in practice in the works of Dimond et al. [101, 102]. In the context of speculative design [25], Martins [245] emphasize the need for intersectional [88, 248, 89] feminist lens in critiquing and dismantling structures of oppression. Feminist theory, methods, and epistemologies have also influenced AI research [4, 5, 378, 67, 350]. Erscoi et al. [121] highlight how women are erased from and by AI technologies. Leavy et al. [223] propose ethical data curation approaches grounded in feminist principles. Using feminist epistemology, Huang et al. [182] critique existing practices of explainable AI, and Varon and Peña [360] critique practices for obtaining digital consent in data extractivist practices in AI. Gender theory have also been employed in these fields in the forms of Queer HCI [232] and Queer AI [209]. Both highlight queering [336, 65] as a tactic to challenge the basis on which categories are constructed.
@@
Irani et al. [186] define postcolonial computing as one that is “centered on the questions of power, authority, legitimacy, participation, and intelligibility in the contexts of cultural encounter, particularly in the context of contemporary globalization. […] It asserts a series of questions and concerns inspired by the conditions of postcoloniality”. Avle et al. [26] criticize the “colonizing impulse” to valorize “universal methods” that are supposedly appropriate across cultural and geopolitical boundaries; instead we can draw from works [15, 387, 386, 320, 131, 10, 11, 310] that center on and incorporate indigenous and non-western values and ethics in the critique and development of technologies. Ali [12] argue for decolonial computing over postcolonial which he criticizes as “Eurocentric critique of Eurocentrism” that “tends to privilege cultural issues over political-economic concerns” and “is noticeably silent about past injustices and does not engage with the matter of reparations”.
@@
Kaba [199] define abolition as “a long-term project and a practice around creating the conditions that would allow for the dismantling of prisons, policing, and surveillance and the creation of new institutions that actually work to keep us safe and are not fundamentally oppressive”. The movement challenges us to move beyond the default assumptions and world views of the carceral state and to dismantle the prison-industrial complex. Incorporating abolitionist values in computing requires us to oppose carceral technologies, surveillance technologies, and military applications [113].
@@
Post-capitalist computing assumes “a socio-economic model that completely replaces capital as the primary method of organising society” [127]. Among other subjects, research in this area contends with, “the racialized dynamics of labor competition” [185], dismantling big tech’s concentration of power [363, 334], and imagining post-work futures [66, 72, 335]. Perhaps, the challenges in this area are best summed up in the words of Jameson [190]: “it is easier to imagine the end of the world than to imagine the end of capitalism”.
@@
These different political lens lend to imagining new futures of computing but there are some themes that cut across them. Firstly, they all recognize technology and society as mutually shaping, and reject both technological determinism [161] and the frame in which technology exists in, what Pfaffenberger [286] calls, a fetishised form [246] where technology is disembodied and disconnected from social relations. Secondly, they recognize that the perspectives, goals, and approaches across this spectrum while sometime distinct are also intersecting. Finally, they all call for structural changes and progress towards alternative futures for society and computing. Perhaps, these aspirations are best articulated by Keyes et al. [208]: “radically reorienting the field towards creating prefigurative counterpower—systems and spaces that exemplify the world we wish to see, as we go about building the revolution in increment”. To affect said changes we need to both recognize the politics of our work and ground it in broader context of political actions [380, 266, 160, 381, 397].

### v4

@@ 2.1 Critical perspectives on fairness and ethics research in IR
Similarly, an important question that all transparency research must contend with is: transparency towards what end? Some works (Polley et al., 2021; Schmitt et al., 2022) motivate transparency as a means to increase user trust in the system. However, we should be critical of whether that trust is warranted, or whether transparency mechanisms could in fact draw users into a false sense of safety and distract them from noticing how the system surveils them and subtly manipulates their behavior (Ravenscraft, 2020; Morrison, 2021). Indeed, Hollanek (2023) argue that “only the sort of transparency that arises from critique—a method of theoretical examination that, by revealing pre-existing power structures, aims to challenge them—can help us produce technological systems that are less deceptive and more just”. It is that kind of critical reflection that we believe should inform transparency research in IR and lead the community towards explicit goals to challenge power (e.g., the power that system owners hold over users) and safeguard user agency.
@@
On the same lines, misinformation research should be motivated by sociotechnical visions for the future of democratic societies, public health, and knowledge production. When we broaden that frame, it becomes apparent that the community must not only focus on automated fact checking, an important research problem, but also understand the social, political, and economic conditions under which misinformation and disinformation is produced and disseminated. The focus of misinformation research then should include identifying, understanding, and addressing the structural mechanisms of misinformation—e.g., data voids (Golebiewski and Boyd, 2019)—as well as ground itself in the articulation of IR’s role in online knowledge production, public health education, and information literacy.
@@
Our perspective here on ongoing research on fairness, transparency, and ethics in IR should not be misconstrued as an argument for doing less of this kind of work. Instead, we believe that the community should be explicit and more ambitious about the changes it wants to affect in broader society and conduct research with a clear mapping between the research goals and the desired social impact. Just as Johnson (2014) challenges the notion that open data directly leads to information justice, we want the IR community to be cautious in their assumption that working on narrowly defined questions of fairness, transparency, and ethics necessarily contributes towards practical social good. To be effective in that endeavour, we believe that we should be explicit in articulating our collective visions for our sociotechnical futures, the changes we want to affect in society, and how we envision our research can bring about those changes.
@@ 3 Towards sociotechnical change
Sociotechnical imaginaries are not born in vacuum. They are moulded and shaped by our values and our politics. Deliberation over what futures we want to bring into being is essentially political, and challenges us to critically reflect on our community’s shared, and yet pluralistic, political values. While the call for explicit political reflection in IR may come as a surprise to some, we need to recognize that our research and the artefacts we produce do not exist outside of the current sociopolitical order (Friedman and Kahn Jr, 2007; Flanagan et al., 2008; Miller, 2021) but as essential cogs in the system, and the absence of political reflection does not imply an absence of politics in our work, but rather translates to implicit complicity in propping up the status quo and neoliberalism (Dourish, 2010; Feltwell et al., 2018; Keyes et al., 2019). Instead, we can learn from how some of our neighboring fields, e.g., HCI and AI, have engaged with these questions, and reflect on how politics shapes and intersects with our own research agendas.
3.1 Prefigurative politics in other IR-adjacent fields
@@
There are several strands of research in IR-adjacent fields that explicate prefigurative politics (Asad, 2019) and ground research in humanistic (Bardzell and Bardzell, 2016a, 2015; Werthner et al., 2024), anti-oppressive and emancipatory (Smyth and Dimond, 2014; Bardzell and Bardzell, 2016a; Kane et al., 2021; Monroe-White, 2021; Saxena et al., 2023), feminist (Wajcman, 2004, 2010; Bardzell, 2010; Bardzell and Bardzell, 2011; Bardzell et al., 2011; Bardzell and Bardzell, 2016b; Bardzell, 2018; D’ignazio and Klein, 2020), queer (Light, 2011; Klipphahn-Karge et al., 2024; Guyan, 2022), postcolonial and decolonial (Irani et al., 2010; Philip et al., 2012; Dourish and Mainwaring, 2012; Sun, 2013; Ali, 2014, 2016; Akama et al., 2016; Irani and Silberman, 2016; Adams, 2021; Mohamed et al., 2020), anti-racist (Abebe et al., 2022), anti-casteist (Kalyanakrishnan et al., 2018; Sambasivan et al., 2021; Vaghela et al., 2022a, b; Shubham, 2022; Kanjilal, 2023), anti-ableist (Williams et al., 2021; Sum et al., 2024), anti-fascist (McQuillan, 2022), abolitionist (Benjamin, 2019; Barabas, 2020; Earl, 2021; Jones and Melo, 2021; Williams and Haring, 2023), post-capitalistic (Feltwell et al., 2018; Browne and Green, 2022), and anarchist (Keyes et al., 2019; Linehan and Kirman, 2014; Asad et al., 2017) epistemologies. Reviewing this full body of literature is out-of-scope of this work but we briefly present a sample to draw from and motivate new IR research agendas for sociotechnical change.
@@
Bardzell and Bardzell (2016a) define humanistic HCI as “any HCI research or practice that deploys humanistic epistemologies (e.g., theories and conceptual systems) and methodologies (e.g., critical analysis of designs, processes, and implementations; historical genealogies; conceptual analysis; emancipatory criticism) in service of HCI processes, theories, methods, agenda-setting, and practices”, and include emanicipatory HCI as an aspiration of humanistic HCI. Kane et al. (2021) propose to incorporate emancipatory pedagogy (Freire, 2020) that does “not advocate the oppressed simply rise and overthrow their oppressors. Instead, […] the oppressors and oppressed create new educational processes that would allow them to work together to create a new type of society that was emancipatory for all”.
@@
In STS, there is a body of work (Longino, 1987; Wajcman, 1991; Hubbard, 2001; Turkle, 2004; Herring et al., 2006; von Hellens et al., 2007; Haraway, 2013; Michelfelder et al., 2017) on gendered inequities caused by technology, and how technology and gender relations mutually shape each other (Wajcman, 2004, 2010). In HCI, Bardzell et al. (Bardzell, 2010; Bardzell and Bardzell, 2011; Bardzell et al., 2011; Bardzell and Bardzell, 2016b; Bardzell, 2018) propose to incorporate feminist theories (Kolmar and Bartkowski, 1999; Friedan, 2010) into research and practice. Bardzell (2010) posits that feminist theories can contribute to interaction design both by critiquing and by generating new insights that inform and shape designs and design processes. We can see feminist HCI in practice in the works of Dimond et al. (Dimond, 2012; Dimond et al., 2013). In the context of speculative design (Auger, 2013), Martins (2014) emphasize the need for intersectional (Crenshaw, 1989; McCall, 2005; Crenshaw, 2013) feminist lens in critiquing and dismantling structures of oppression. Feminist theory, methods, and epistemologies have also influenced AI research (Adam, 1995, 2013; Wellner and Rothman, 2020; Browne et al., 2023; Toupin, 2024). Erscoi et al. (2023) highlight how women are erased from and by AI technologies. Leavy et al. (2021) propose ethical data curation approaches grounded in feminist principles. Using feminist epistemology, Huang et al. (2022) critique existing practices of explainable AI, and Varon and Peña (2021) critique practices for obtaining digital consent in data extractivist practices in AI. Gender theory have also been employed in these fields in the forms of Queer HCI (Light, 2011) and Queer AI (Klipphahn-Karge et al., 2024). Both highlight queering (Sta, 1997; Brooks et al., 2021) as a tactic to challenge the basis on which categories are constructed.
@@
Irani et al. (2010) define postcolonial computing as one that is “centered on the questions of power, authority, legitimacy, participation, and intelligibility in the contexts of cultural encounter, particularly in the context of contemporary globalization. […] It asserts a series of questions and concerns inspired by the conditions of postcoloniality”. Avle et al. (2017) criticize the “colonizing impulse” to valorize “universal methods” that are supposedly appropriate across cultural and geopolitical boundaries; instead we can draw from works (Alsheikh et al., 2011; Wong, 2012; Winschiers-Theophilus and Bidwell, 2013; Shaw et al., 2014; Fox and Le Dantec, 2014; Ahmed et al., 2015; Akama et al., 2016; Sambasivan et al., 2021) that center on and incorporate indigenous and non-western values and ethics in the critique and development of technologies. Ali (2014) argue for decolonial computing over postcolonial which he criticizes as “Eurocentric critique of Eurocentrism” that “tends to privilege cultural issues over political-economic concerns” and “is noticeably silent about past injustices and does not engage with the matter of reparations”.
@@
Kaba (2021) define abolition as “a long-term project and a practice around creating the conditions that would allow for the dismantling of prisons, policing, and surveillance and the creation of new institutions that actually work to keep us safe and are not fundamentally oppressive”. The movement challenges us to move beyond the default assumptions and world views of the carceral state and to dismantle the prison-industrial complex. Incorporating abolitionist values in computing requires us to oppose carceral technologies, surveillance technologies, and military applications (Earl, 2021).
@@
Post-capitalist computing assumes “a socio-economic model that completely replaces capital as the primary method of organising society” (Feltwell et al., 2018). Among other subjects, research in this area contends with, “the racialized dynamics of labor competition” (Irani, 2018), dismantling Big Tech’s concentration of power (Verdegem, 2022; Srnicek, 2017), and imagining post-work futures (Browne and Green, 2022; Butler, 2018; Srnicek and Williams, 2015). Perhaps, the challenges in this area are best summed up in the words of Jameson (2003): “it is easier to imagine the end of the world than to imagine the end of capitalism”.
@@
These different political lens lend to imagining new futures of computing but there are some themes that cut across them. Firstly, they all recognize technology and society as mutually shaping, and reject both technological determinism (Greene et al., 2019) and the frame in which technology exists in, what Pfaffenberger (1988) calls, a fetishised form (Marx, 1867) where technology is disembodied and disconnected from social relations. Secondly, they recognize that the perspectives, goals, and approaches across this spectrum while sometime distinct are also intersecting. Finally, they all call for structural changes and progress towards alternative futures for society and computing. Perhaps, these aspirations are best articulated by Keyes et al. (2019): “radically reorienting the field towards creating prefigurative counterpower—systems and spaces that exemplify the world we wish to see, as we go about building the revolution in increment”. To affect said changes we need to both recognize the politics of our work and ground it in broader context of political actions (Wickenden, 2018; Moore, 2020; Green, 2021; Widder et al., 2023; Young et al., 2021).
3.2 Proposals for IR